How much of a 50% acid solution must be mixed with 80 gallons of a 14% acid solution to obtain a solution that is 40% acid?
I feel like I've done something like this before, but I just can't get started. Thank you for any help!

Answer by Earlsdon(6294)

Let x = the required number of gallons of the 50% acid solution.
Write the equation in terms of the amount of acid in the solution:
For example: 80 gallons of 14% acid solution would have 0.14(80) = 11.2 gallons of acid. So...
0.5x+0.14(80) = 0.4(80+x) Simplify and solve for x.
0.5x+11.2 = 32+0.4x Subtract 0.4x from both sides.
0.1x+11.2 = 32 Subtract 11.2 from both sides.
0.1x = 20.8 Finally, divide both sides by 0.1
x = 208 gallons of the 50% acid solution.